OpenStudy (anonymous):

Given f(x) = 5x+6 and g(x) = 5x^2 find ) (f+g)(x)

geerky42 (geerky42):

\((f+g)(x)\text{ is same as }f(x)+g(x)\)

geerky42 (geerky42):

So \((f+g)(x) = f(x)+g(x) = [5x+6]+[5x^2]\)

OpenStudy (anonymous):

can you tell me how you would find the domain of this equation?

geerky42 (geerky42):

Ok so \(5x^2+5x+6\) is polynomial (trinomial). So domain is \((-\infty,\infty)\) or \(\mathbb{R}\)

geerky42 (geerky42):

Because there is no such input that would lead output to be undefined